OpenStudy (anonymous):

Find the area of the region bounded by the curves y = x2 and y = cos(x). Give your answer correct to 2 decimal places

OpenStudy (perl):

y = x^2 and y = cos x

OpenStudy (perl):

Is that what you meant?

OpenStudy (freckles):

A long long time ago x2 use to mean x^2. But nowadays I only see its interpretation as 2 times x.

OpenStudy (perl):

Yes , otherwise they would write 2x (generally speaking).

OpenStudy (perl):

not x2

OpenStudy (perl):

@Krissy3039 Please show any work you have done.

OpenStudy (amistre64):

to avoid confusion i still write xx :)

OpenStudy (freckles):

what about for x^(1/2)

OpenStudy (amistre64):

X > maybe?

OpenStudy (amistre64):

sqrt(x) should work tho

OpenStudy (tkhunny):

Learn some LaTeX!

OpenStudy (freckles):

agreed :p

OpenStudy (perl):

$$ \Large y = x^2 , ~y = \cos x $$

OpenStudy (tkhunny):

Okay, to the integral!!!

OpenStudy (amistre64):

latex warps the brain ... you start writing by hand and the latex just doesnt look the same anymore.

OpenStudy (tkhunny):

...said the person who probably has legible handwriting. :-) LaTeX has connected me with the world! People can read what I write!

OpenStudy (perl):

It helps to graph the two curves to find the region that is bounded. The region where the two areas overlaps is going to be the area that we want. https://www.desmos.com/calculator/xfnjcxr8vq

OpenStudy (anonymous):

it was x^2 sorry

OpenStudy (freckles):

looking at @perl 's graph you have two intersections to find . The solutions can be found by solving x^2=cos(x). There is no algebraic way to solve that equation though. You will have to approximate those intersections.

OpenStudy (tkhunny):

You can get in the neighborhood with a quadratic series approximation of the cosine. \(1 - \dfrac{x^{2}}{2} = x^{2}\implies x = 0.816\) or so With just a hair more difficulty, still only an algebra problem, you can use the quartic approximation. \(1 - \dfrac{x^{2}}{2} + \dfrac{x^{4}}{24} = x^{2} \implies x = 0.824\) That's actually pretty close.

OpenStudy (anonymous):

thats the x value for the intercept and the y value is .679

OpenStudy (anonymous):

what would you do to find the area then

OpenStudy (tkhunny):

Of course, I meant, \(Quadratic \implies x = \dfrac{\sqrt{6}}{3}\;and\;Quartic \implies x = \sqrt{18-10\sqrt{3}}\). What would be the point of the approximation if the answer weren't prettier? Where is your integral with the limits in place?

OpenStudy (tkhunny):

Those are only approximations. You will need to do better than that. That just gives you a place to start looking. The first is too far left and the second too far right. You can bracket the actual answer with that sort of information. Let's see that integral.

OpenStudy (anonymous):

wait what im confused

OpenStudy (freckles):

do you know how to setup the integral in that region where there is purple and green in @perl 's little page he posted above as a link

OpenStudy (freckles):

\[\int\limits_{\text{ \lower limit (the left intersection x value of course )}}^{\text{ upper limit (the right intersection x value of course )}}(\text{ top function } - \text{ bottom function } ) dx\]

OpenStudy (anonymous):

okay so would it be x^3/3 - sin(x)?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

on the interval in which we look at cos(x) and x^2 isn't cos(x)>x^2?

OpenStudy (freckles):

so it should be the other way around sin(x)-x^3/3

OpenStudy (freckles):

now you just have to find your intersection and do the difference of the expression above evaluated at both

OpenStudy (anonymous):

i got - 1.46787...

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

what did you use as your interesection numbers?

OpenStudy (anonymous):

-.8241 and .8241

OpenStudy (freckles):

hmm... so you did: (sin(.8241)-(.8241)^3/3)-(sin(-.8241)-(-.8241)^3/3)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then would the area be 1.09?

OpenStudy (freckles):

\[(\sin(.8241)-\frac{(.8241)^3}{3})-(\sin(-.8241)-\frac{(-.8241)^3}{3})\] maybe you didn't exactly enter it into the calculator right \[\sin(.8241)-\frac{(.8241)^3}{3}-(-\sin(.8241)-\frac{-(.8241)^3}{3}) \\ \text{ since } \sin(x) \text{ and } x^3 \text{ are odd functions } \\ 2 \sin(.8241)-2 \cdot \frac{(.8241)^3}{3} \\ =2(\sin(.8241)-\frac{2}{3}\sin(.8241))\] yeah that sounds a lot closer

OpenStudy (anonymous):

so should i go with 1.09 then?

OpenStudy (freckles):

it is a pretty good approximation

OpenStudy (anonymous):

okay thank you!

OpenStudy (perl):

\(\color{blue}{\text{Originally Posted by}}\) @perl \[ \Large{ \int_{-.824}^{.824} \cos (x)-x^2 ~dx ~ \\~\\~\\ or \\~\\ 2\cdot \int_{0}^{.824} \cos (x) -x^2 ~dx ~ =1.0947 }\] \(\color{blue}{\text{End of Quote}}\)

OpenStudy (freckles):

and i made a type-o above

OpenStudy (freckles):

\[\sin(.8241)-\frac{(.8241)^3}{3}-(-\sin(.8241)-\frac{-(.8241)^3}{3}) \\ \text{ since } \sin(x) \text{ and } x^3 \text{ are odd functions } \\ 2 \sin(.8241)-2 \cdot \frac{(.8241)^3}{3} \\ =2(\sin(.8241)-\frac{1}{3}(.8241)^3)\] since I dragged that 2 out of there